OpenStudy (anonymous):

Solve the proportion below for x. x/18=3/12

OpenStudy (anonymous):

cross multiply.

OpenStudy (anonymous):

@bggjdiewj I know that's what I'm supposed to do..... I just don't know how to cross multiply haha I learned it in like 8th grade but I forget now :/

OpenStudy (anonymous):

Like 18 by 3 and x by 12. Then your last fraction you simply as much as you can.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

Kind of, could I use 1 instead of x? or would that mess it up?

OpenStudy (anonymous):

@bggjdiewj so far I have 54/12

OpenStudy (raden):

yes, x = 54/12 = ... just simplify then

OpenStudy (anonymous):

@RadEn 9/2?

OpenStudy (raden):

nice, that's right!

OpenStudy (anonymous):

Thank you! :)

OpenStudy (raden):

welcome :)

OpenStudy (anonymous):

@RadEn wait, what would x be though? I'm still confused :(

OpenStudy (raden):

u got x = 9/2, right ?

OpenStudy (anonymous):

yeah

OpenStudy (raden):

so, what the trouble then ...

OpenStudy (anonymous):

Oh wait, so x = 9/2 meaning 9/2 is my final answer? I feel extremely stupid now.... Nevermind haha, I'm sorry! I've been working on this all day and I'm really tired

OpenStudy (raden):

yes, that's the final answer :)